OpenStudy (anonymous):

What is the solution set for |2x – 6| – 7 = 7? Answer x = –4 and x = –10 x = 4 and x = 10 x = –4 and x = 10 x = 4 and x = –10

OpenStudy (aroub):

x=-4 x=10

OpenStudy (anonymous):

x=-4 x=10

OpenStudy (anonymous):

|2x – 6| – 7 = 7 |2x – 6| = 14 2x – 6 = 14 2x – 6 = -14 x= (14+6)/2 x = (-14+6)/2 x = 10 x = - 4